W. C. WoodW. C. Wood was the largest freezer manufacturer in Canada. It had two plants in Guelph, Ontario; one in the United States; and another in Mexico. In the early 2000s, some appliance manufacturers, such as Whirlpool, had converted from using cardboard box packaging to using plastic stretch wrap with corner padding (called clear-view packaging). The VP of manufacturing was wondering whether to also convert. The damage warranty rate for cardboard box packaging (1–3 percent depending on the type of freezer) was expected to remain the same with clear-view packaging. For simplicity, we consider only one of the Guelph plants with one line. Whereas a cardboard box required two workers to box an appliance, the clear-view packaging would be automated. Labour and material costs would be reduced by $0.50 and $1.50 per unit, respectively. However, electricitycosts would increase by $10,000 per year. The stretch wrap equipment by MSK Covertech cost $130,000. The fitting of corner posts on each freezer required another machine, which cost $140,000. In addition, all lift truck arms had to be converted to clamp type at a total cost of $30,000.Questionsa.What is the total variable cost saving per freezer?b.What is the total investment cost per year if all pieces of equipment are expected to last 10 years (use straight-line depreciation)?c.How many freezers per year should the Guelph plant produce in order to break even in the clear-view packaging conversion?
Answer #1

a. Calculation of Total Variable Cost saving per freezer unit :-

Particulars Amount (In $)
Labour Cost Reduced per unit 0.50
Material Cost Reduced per unit 1.50
Total Cost Reduced per unit 2.00

Thus, the total variable cost saving per freezer unit by using plastic stretch wrap instead of using cardboard box packaging is $ 2 per freezer unit.

b. Calculation of Total Investment Cost per year :-

1 Total Cost of Equipment

Particulars Amount (In $)
Stretch Wrap equipment    1,30,000.00
Machine for fitting corner posts    1,40,000.00
Cost of conversion of lift truck arms to clamp type       30,000.00
Total Cost of Equipment    3,00,000.00

It is given in the question that all pieces of equipment are expected to last 10 years and Straight line depreciation is to be followed. Thus, depreciation cost for all equipment is as follows :-

\textrm{Depreciation under Straight Line Method =}\frac{Cost - Residual Value}{Life of Asset}

In the given question, the cost of all equipment is $ 3,00,000.

Residual Value of equipments at the end of useful life of asset is not given, thus we assume it to be Nil.

Life of Asset is 10 years. Thus,

\textrm{Depreciation under Straight Line Method =}\frac{300,000-0}{10}

Depreciation under Straight Line Method = $ 30,000 per year.

Thus, calculation of total investment cost per year is :-

Particulars Amount (In $)
Electricity Cost       10,000.00
Depreciation cost of all equipment       30,000.00
Total Investment cost per year       40,000.00

Thus, the investment cost per year is $ 40,000 for using plastic stretch wrap instead of using cardboard box packaging.

c. Calculation of breakeven point :-

\textrm{Breakeven point in number of units =}\frac{Total Cost}{Savings per unit}

\textrm{Breakeven point in number of units =}\frac{40,000}{2}

Breakeven point = 20,000 units.

Thus, Guelph plant should produce 20,000 freezers per year in order to breakeven in the clear-view packaging conversion.